OpenStudy (anonymous):

Binary subtraction 111 - 11001

OpenStudy (anonymous):

I've used r's complement method but I get the wrong answer: Two's complement: 00111 Adding them: 111+111=1110 Omitting carry: 110 And that should be my answer, except it's incorrect.

OpenStudy (anonymous):

I have no idea how to do this stuff, so I just looked at an online binary calculator for you, and the result was this -10010 I've done binary addition before, but never this

OpenStudy (anonymous):

That's the answer I'm trying to get to, the method above worked on every problem so far. But got stuck on this one.

OpenStudy (anonymous):

Well, I have absolutely NO idea how to give an explanation for that, sorry :/

OpenStudy (amistre64):

add backwards :) 11001 - 00111 ------- 0 0+1=1, carry a 0 0 11001 - 00111 ------- 10 1+1=0, carry a 1 10 11001 - 00111 ------- 010 etc ... 110 11001 - 00111 ------- 0010 0110 11001 - 00111 ------- 10010 0110

OpenStudy (amistre64):

111 - 11001 can be written in addition by 2comping the negative, swap out and add 1 111 + 00110 + 1 now add 00111 00111 ------ 01110 convert it back, swap and add 1 10001 + 1 = 10010

OpenStudy (anonymous):

That's brilliant Thanks for the alternative way, my brain was frying over the first one.